The volume of a sphere is to be computed from the measured value of its radius. Estimate the maximum permissible percentage error in the measurement if the percentage error in the volume must be kept within ±6%.


1
Expert's answer
2021-10-01T04:22:23-0400

Let V be the volume of the sphere and r be its radius.

V = (4/3)π r 3

Δ ​V/V = (3(4/3) π r2r/((4/3π r3)

= 3Δ r/r

Δ ​V/V = ±6%

Δ r/r = ±2%

 

Answer: Maximum permissible error in the measurement of

its radius is ± 2%